# Generalized Pulse Design in Floquet Engineering: Application to Interacting Spin Systems

### v1

@@ V Summary and Outlook
We constructed a formalism that loosens the set of assumptions [23] underlying pulse shapes for Floquet engineering of interacting spin systems into accurate effective Hamiltonians. Our formalism allows tunability of pulse shape, e.g., width and strength, while maintaining controlled error. Specific examples using global Pauli-$x$ and Pauli-$z$ pulses were implemented to demonstrate conversion of long-range XXZ spin models to Ising models. The formalism constructed here is accurate to lowest order in inverse frequency of the driving pulses, $\mathcal{O}(\omega^{0})$.
@@
It is also straightforward to adapt the formalism discussed here to models containing any operators respecting the Lie algebra commutation relations, not just $SU(2)$ spin systems. Future work will allow Floquet-engineering of more general multi-component $SO(N)$ models relevant to rotational states of polar molecules (See Appendix D) or $SU(N)$ models relevant to other states of molecules [45, 46]. In such cases, $F^{\alpha}_{j}[\hat{H}]$ will be modified to include new structure constants in commutation relations.
@@ Appendix A Proof of Generalization
In this section we prove that we generalize the family of pulses defined by Eq. K1 in Ref. 23:
@@ Appendix B Full Derivation of the Lowest Order Hamiltonian and Corrections
where $NT$ is the cycle time. Now consider the gauge transformation in Eq. (3), which gives rise to the effective Schrödinger equation: $\hat{H}_{\text{eff}}\phi=i\partial_{t}\phi$. Furthermore, if the effective timescale of interaction of the Hamiltonian $H_{0}$ is much larger than the timescale set by $\omega^{-1}$, we can use Eqs. (15) and (16). We substitute Eqs. (15) and  (16) into Eqs. (3) and use the identities:
$e^{i\hat{K}(t)}\hat{H}e^{-i\hat{K}(t)}=\sum_{p=0}^{\infty}\frac{i^{p}}{p!}[\hat{K},...[\hat{K},\hat{H}]...]$ (46)
$\left(\frac{\partial e^{i\hat{K}(t)}}{\partial t}\right)e^{-i\hat{K}(t)}=\sum_{p=1}^{\infty}\frac{i^{p}}{p!}\left[\hat{K},...\left[\hat{K},\frac{\partial\hat{K}}{\partial t}\right]...\right],$ (47)
where the commutators occur $p$ times in each of the sums. The substitution gives rise to the full expansion:
$+\omega\sum_{\alpha}\sum_{j}g^{\alpha}_{j}(t)\left\{\sum_{p=0}^{\infty}\frac{i^{p}}{p!}\sum_{n_{1},...,n_{p}=0}^{\infty}\omega^{-(n_{1}+...+n_{p})}[\hat{K}^{(n_{1})},...[\hat{K}^{(n_{p})},\hat{\sigma}^{(\alpha)}_{j}]...]\right\}$
$+i\sum_{q=0}^{\infty}\sum_{p=1}^{\infty}\omega^{-q}\frac{i^{p}}{p!}\left\{\sum_{n_{2},...,n_{p}=0}^{\infty}\omega^{-(n_{2}+...+n_{p})}\left[\hat{K}^{(n_{2})},...\left[\hat{K}^{(n_{p})},\frac{\partial\hat{K}^{(q)}}{\partial t}\right]...\right]\right\}.$ (48)
The lowest order term occurring on the left-hand side is $\mathcal{O}(\omega^{0})$, but on the right-hand side the second and third term of the expansion are $\mathcal{O}(\omega)$. Since $\partial\hat{K}/\partial t\approx\omega\hat{K}$, we infer that $\hat{K}^{(0)}$ must satisfy:
$\omega\sum_{p=0}^{\infty}\frac{i^{p}}{p!}\left[\hat{K}^{(0)},...\left[\hat{K}^{(0)},\sum_{\alpha}\sum_{j}\hat{\sigma}^{(\alpha)}_{j}g^{\alpha}_{j}(t)\right]...\right]$ (49)
$+i\sum_{p=1}^{\infty}\frac{i^{p}}{p!}\left[\hat{K}^{(0)},...\left[\hat{K}^{(0)},\frac{\partial\hat{K}^{(0)}}{\partial t}\right]...\right]=0.$
where the commutator occurs $p$ times in the first term and $p-1$ times in the second term. If $\hat{K}^{(0)}$ commutes with its time derivative, this simplifies to Eq. (17) with solutions given by Eq. (18).
Equation (19) has a gauge freedom in the lower bound which can be selected according to what is convenient [24]. The $G^{\alpha}_{j}$ are periodic of period $NT$ since the $g^{\alpha}_{j}(t)$ are periodic and cyclic of period $NT$. It is also convenient to impose the cyclic condition on the $G^{\alpha}_{j}$ as well. Neglecting all higher-order contributions, these observations together imply:
$\hat{H}^{(0)}_{\text{eff}}$ $=\sum_{p=0}^{\infty}\frac{i^{p}}{p!}\sum_{\alpha_{1},...,\alpha_{p}}\sum_{j_{1},...,j_{p}}G^{\alpha_{1}}_{j_{1}}\cdots G^{\alpha_{p}}_{j_{p}}[\hat{\sigma}^{(\alpha_{1})}_{j_{1}},...[\hat{\sigma}^{(\alpha_{p})}_{j_{p}},\hat{H}_{0}]...]$
$+\omega\sum_{\alpha}\sum_{j}g^{\alpha}_{j}(t)\bigg{\{}\sum_{p=0}^{\infty}\frac{i^{p}}{p!}\sum_{\alpha_{1},...,\alpha_{p}}\sum_{j_{1},...,j_{p}}G^{\alpha_{1}}_{j_{1}}\cdots G^{\alpha_{p}}_{j_{p}}[\hat{\sigma}^{(\alpha_{1})}_{j_{1}},...[\hat{\sigma}^{(\alpha_{p})}_{j_{p}},\hat{\sigma}^{(\alpha)}_{j}]...]\bigg{\}}+\mathcal{O}(\omega^{-1}).$ (50)
Using the assumptions of Sec. III.1.3, it follows that the $G^{\alpha}_{j}$ have the form:
@@
Furthermore, it is convenient (but not necessary) to assume the $g^{\alpha}_{j}(t)$ are orthogonal to the $G^{\alpha}_{j}$ in the sense that
@@
which can be enforced, for example, by choosing $g(t)=v\cos(\omega t)$. Lastly, we can take the time average of both sides in Eq. (50) to obtain Eq. (20) with the commutator of order $2p$ and when the commutator at $p=0$ is identified with the identity operator on $\hat{H}_{0}$. We have therefore derived the $\mathcal{O}(\omega^{0})$ terms in Eq. (20).
To obtain the general structure of $\mathcal{O}(\omega^{-1})$ corrections, we follow the same line of reasoning, identifying the differential equation that $\hat{K}^{(1)}$ must satisfy, and equating orders of powers of $\omega^{-1}$. If we carry out the expansion to the next highest order in $\omega^{-1}$, we obtain the general differential equation:
@@
$\begin{split}&\omega\sum_{p=1}^{\infty}\frac{i^{p}}{p!}D^{p}_{\{0,1\}}\left[\sum_{\alpha}\sum_{j}\hat{\sigma}^{(\alpha)}_{j}g^{\alpha}_{j}(t)\right]\\ &+\sum_{p=1}^{\infty}\frac{i^{p}}{p!}\left(D^{p-1}_{\{0\}}\left[\frac{\partial\hat{K}^{(1)}}{\partial t}\right]+D^{p}_{\{0,1\}}\left[\frac{\partial\hat{K}^{(0)}}{\partial t}\right]\right)=0,\end{split}$ (53)
where $D^{p}_{\{0,1\}}[-]$ is a sum over all permutations of a single appearance of $\hat{K}^{(1)}$ and $p-1$ appearances of $\hat{K}^{(0)}$ in the Lie derivative. If we assume that we solve Eq. (16) and keep only the lowest order terms here, we obtain another simplified differential equation for $\hat{K}^{(1)}$:
$i\omega\sum_{j,\alpha}g^{\alpha}_{j}(t)[\hat{K}^{(1)},\hat{\sigma}^{(\alpha)}_{j}]-\frac{\partial\hat{K}^{(1)}}{\partial t}=0,$ (54)
which is reminiscent of the Heisenberg equation of motion. Notice that even if we keep higher order terms in these, both yield a differential equation which can be solved for $\hat{K}^{(0)}$, $\hat{K}^{(1)}$ in terms of $\hat{V}(t)$, and in such a way that they are not coupled together. This means that in principle the corrections due to neglecting these terms can be worked around within the expansion. This implies, when solved, that $\hat{K}^{(1)}$ is a function of $\hat{V}(t)$, and we have our next order correction:
$\hat{H}^{(1)}=\sum_{p=1}^{\infty}\frac{i^{p}}{p!}D^{p}_{\{0,1\}}[\hat{H}_{0}]+\mathcal{O}\left(\hat{K}^{(2)},\omega^{-2}\right),$ (55)
after neglecting all terms that either involve $\hat{K}^{(2)}$ or higher powers of $\omega^{-1}$. When applied to XXZ-model examples, we find that the correction goes like:
$\hat{H}_{\text{eff}}^{\text{S}}\sim\hat{H}^{\text{S},(0)}_{\text{eff}}+\mathcal{O}\left(\frac{\max[|J_{\perp}|,|J_{z}|]}{\omega}\right).$ (56)
This shows that corrections to terms discussed in the main text are of order $\omega^{-1}$.
@@ Appendix C Full Derivation of the Ising Interaction from the XXZ Hamiltonian
$\hat{H}^{\text{S},(0)}_{\text{eff}}$ $=\sum_{p=0}^{\infty}\frac{(-1)^{p}v^{2p}}{(2p)!}\overline{G_{1}^{2p}}\left[\sum_{k_{1}}\hat{\sigma}^{(1)}_{k_{1}},...\left[\sum_{k_{2p}}\hat{\sigma}^{(1)}_{k_{2p}},\hat{H}_{0}^{\text{S}}\right]...\right]$
$+\sum_{p=0}^{\infty}\frac{(-1)^{p}v^{2p}}{(2p)!}\overline{G_{2}^{2p}}\left[\sum_{k_{1}}\hat{\sigma}^{(2)}_{k_{1}},...\left[\sum_{k_{2p}}\hat{\sigma}^{(2)}_{k_{2p}},\hat{H}_{0}^{\text{S}}\right]...\right].$
@@
$D^{2p}_{1}[\hat{H}_{XY}]$ $:=\sum_{k_{1}}\cdots\sum_{k_{2p}}\left[\hat{\sigma}^{(1)}_{k_{1}},...\left[\hat{\sigma}^{(1)}_{k_{2p}},\hat{H}_{XY}\right]...\right]$ (57)
$D^{2p}_{1}[\hat{H}_{ZZ}]$ $:=\sum_{k_{1}}\cdots\sum_{k_{2p}}\left[\hat{\sigma}^{(1)}_{k_{1}},...\left[\hat{\sigma}^{(1)}_{k_{2p}},\hat{H}_{ZZ}\right]...\right]$ (58)
$D^{2p}_{2}[\hat{H}_{XY}]$ $:=\sum_{k_{1}}\cdots\sum_{k_{2p}}\left[\hat{\sigma}^{(2)}_{k_{1}},...\left[\hat{\sigma}^{(2)}_{k_{2p}},\hat{H}_{XY}\right]...\right]$ (59)
$D^{2p}_{2}[\hat{H}_{ZZ}]$ $:=\sum_{k_{1}}\cdots\sum_{k_{2p}}\left[\hat{\sigma}^{(2)}_{k_{1}},...\left[\hat{\sigma}^{(2)}_{k_{2p}},\hat{H}_{ZZ}\right]...\right].$ (60)
@@
$+\sum_{k,k^{\prime}}D^{1}_{k,k^{\prime}}[\hat{\sigma}^{(2)}_{j}\hat{\sigma}^{(2)}_{j^{\prime}}]\bigg{)},$
@@
$D^{2}[\hat{H}_{0}^{\text{S}}]$ $=\frac{J_{\perp}}{2}\sum_{j\neq j^{\prime}}\mathcal{V}_{j,j^{\prime}}\sum_{k,k^{\prime}}\left(D^{1}_{k,k^{\prime}}[\hat{\sigma}^{(2)}_{j}\hat{\sigma}^{(2)}_{j^{\prime}}]+D^{2}_{k,k^{\prime}}[\hat{\sigma}^{(1)}_{j}\hat{\sigma}^{(1)}_{j^{\prime}}]\right)+\frac{J_{z}}{2}\sum_{j\neq j^{\prime}}\mathcal{V}_{j,j^{\prime}}\sum_{k,k^{\prime}}\left(D^{1}_{k,k^{\prime}}[\hat{\sigma}^{(3)}_{j}\hat{\sigma}^{(3)}_{j^{\prime}}]+D^{2}_{k,k^{\prime}}[\hat{\sigma}^{(3)}_{j}\hat{\sigma}^{(3)}_{j^{\prime}}]\right).$ (61)
@@
$\hat{H}_{\text{eff}}^{\text{S},(0)}$ $=\hat{H}_{0}^{\text{S}}+\frac{1}{16}\left(\sum_{p=1}^{\infty}\frac{(-1)^{p}(4v)^{2p}}{(2p)!}\overline{G^{2p}_{1}}\right)D^{2}_{1}[\hat{H}_{0}^{\text{S}}]$
@@
$\hat{H}^{\text{S},(0)}_{\text{eff}}=\sum_{j\neq j^{\prime}}\mathcal{V}_{j,j^{\prime}}\left[(J_{\perp}/2-4\delta JU^{(2)})\hat{\sigma}^{(1)}_{j}\hat{\sigma}^{(1)}_{j^{\prime}}+(J_{\perp}/2-4\delta JU^{(1)})\hat{\sigma}^{(2)}_{j}\hat{\sigma}^{(2)}_{j^{\prime}}\right]+\sum_{j\neq j^{\prime}}\mathcal{V}_{j,j^{\prime}}\left[\frac{J_{z}}{2}+4\delta J(U^{(1)}+U^{(2)})\right]\hat{\sigma}^{(3)}_{j}\hat{\sigma}^{(3)}_{j^{\prime}}.$ (69)
$\hat{H}^{\text{S},(0)}_{\text{eff}}$ allows us to use Floquet engineering to inter-relate a large class of starting Hamiltonians and effective Hamiltonians. For example, if we want to obtain $\hat{H}^{(0)}_{\text{eff}}\propto\hat{H}_{ZZ}$, we need:
@@
$\hat{H}^{\text{S},(0)}_{\text{eff}}=\frac{J_{z}+2J_{\perp}}{2}\sum_{j\neq j^{\prime}}\mathcal{V}_{j,j^{\prime}}\hat{\sigma}^{(3)}_{j}\hat{\sigma}^{(3)}_{j^{\prime}}.$ (72)

### v2

@@ V Summary and Outlook
We have constructed a formalism that loosens the set of assumptions [goldman2014periodically] underlying pulse shapes for Floquet engineering of interacting spin systems into accurate effective Hamiltonians. Our formalism allows tunability of pulse shape, e.g., width and strength, while maintaining controlled error. Specific examples using global Pauli-$x$ and Pauli-$z$ pulses were implemented to demonstrate conversion of long-range XXZ spin models to Ising models. The formalism constructed here is accurate to lowest order in inverse frequency of the driving pulses, $\mathcal{O}(\omega^{0})$. It would be interesting to compare our Floquet engineering formalism to digital Trotter-based methods (See, e.g., Refs. [PASTORI2022, CHERTKOV2024, ECKSTEIN2024, HAGHSHENAS2025]).
@@
It is also straightforward to adapt the formalism discussed here to models containing any operators respecting the Lie algebra commutation relations, not just $SU(2)$ spin systems. Future work will allow Floquet-engineering of more general multi-component $SO(N)$ models relevant to rotational states of polar molecules (See Appendix D) or $SU(N)$ models relevant to other states of molecules [HOMEIER2024, MUKHERJEE2025] or superconducting qudits [Kiktenko2025]. In such cases, $F^{\alpha}_{j}[\hat{H}$] will be modified to include new structure constants in commutation relations. Interesting models that could be pursued include the Affleck–Kennedy–Lieb–Tasaki [Affleck1987a] models which have interesting valence-bond solid ground states.
Acknowledgements.
We acknowledge support from AFOSR (FA9550-19-1-0272). BG was supported by the AFOSR MURI program under agreement number FA9550-221-0339. RS and VS acknowledge support from AFOSR (FA2386-21-1-4081, and FA9550-23-1-0034) and ARO (W911NF2210247). We thank B. DeMarco and S. Kotochigova for valuable conversations.
@@ Appendix A Proof of Generalization
In this section we prove that we generalize the family of pulses defined by Eq. K1 in Ref. goldman2014periodically:
@@ Appendix B Full Derivation of the Lowest Order Hamiltonian and Corrections
$\int_{0}^{NT}g^{\alpha}_{j}(t)dt=0,$ (45)
where $NT$ is the cycle time. Now consider the gauge transformation in Eq. (3), which gives rise to the effective Schrödinger equation: $\hat{H}_{\text{eff}}\phi=i\partial_{t}\phi$. Furthermore, if the effective timescale of interaction of the Hamiltonian $\hat{H}_{0}$ is much larger than the timescale set by $\omega^{-1}$, we can use Eqs. (17) and (18). We substitute Eqs. (17) and  (18) into Eqs. (3) and use the identities:
$e^{i\hat{K}(t)}\hat{H}e^{-i\hat{K}(t)}$ $=\sum_{p=0}^{\infty}\frac{i^{p}}{p!}[[\hat{K},\hat{H}]]_{p}$ (46)
$\left(\frac{\partial e^{i\hat{K}(t)}}{\partial t}\right)e^{-i\hat{K}(t)}$ $=\sum_{p=0}^{\infty}\frac{i^{p+1}}{(p+1)!}\left[\left[\hat{K},\frac{\partial\hat{K}}{\partial t}\right]\right]_{p},$ (47)
The substitution gives rise to the full expansion:
@@
In each sum, the commutator occurs $p$ times. Assume the lowest order term in the expansion of the kick operator satisfies:
$\hat{K}^{(0)}(t)=\sum_{\alpha,j}G^{\alpha}_{j}(t)\hat{\sigma}^{(\alpha)}_{j}$, (49)
where:
$G^{\alpha}_{j}(t)=\omega\int^{t}dt^{\prime}g^{\alpha}_{j}(t^{\prime}).$ (50)
This follows by requiring that the terms of order $\mathcal{O}(\omega)$ sum to zero on the right-hand side of the full expansion for $H_{\text{eff}}$ (Compare Appendix K in Ref. [goldman2014periodically]).
Returning to Eq. (48), the lowest order of the effective Hamiltonian becomes:
$\hat{H}^{(0)}_{\text{eff}}=\hat{H}_{0}+\sum_{p=1}^{\infty}\frac{i^{p}}{p!}[[\hat{K}^{(0)},\hat{H}_{0}]]_{p}+\hat{R}$, (51)
where $\hat{R}$ is a remainder term that depends on both $\hat{K}^{(0)}$ and $\hat{K}^{(1)}$. In particular,
@@
$+\frac{i}{\omega}\sum_{p=1}^{\infty}\frac{i^{p+1}}{(p+1)!}\sum_{n_{1},...,n_{p}=0}^{1}\delta_{\sum_{j=1}^{p}n_{j},1}\left[\hat{K}^{(n_{1})},...\left[\hat{K}^{(n_{p})},\frac{\partial\hat{K}^{(0)}}{\partial t}\right]...\right]+\frac{i}{\omega}\sum_{p=0}^{\infty}\frac{i^{p+1}}{(p+1)!}\left[\left[\hat{K}^{(0)},\frac{\partial\hat{K}^{(1)}}{\partial t}\right]\right]_{p}$
@@
$=\sum_{p=1}^{\infty}\frac{i^{p}}{p!}\frac{p}{p+1}\sum_{n_{1},...,n_{p}=0}^{1}\delta_{\sum_{j=1}^{p}n_{j},1}\left[\hat{K}^{(n_{1})},...\left[\hat{K}^{(n_{p})},\sum_{\alpha,j}g^{\alpha}_{j}(t)\hat{\sigma}^{(\alpha)}_{j}\right]....\right]+\frac{i}{\omega}\sum_{p=0}^{\infty}\frac{i^{p+1}}{(p+1)!}\left[\left[\hat{K}^{(0)},\frac{\partial\hat{K}^{(1)}}{\partial t}\right]\right]_{p}$ (52)
where $\delta_{\sum_{j}n_{j},1}$ means when the sum of the set $\{n_{j}\}$ is one, the delta function is one, and is zero otherwise. The commutator occurs $p$ times in each term.
To prove Eq. (20), the remainder term (or its time average) must vanish. Importantly, the strong driving forces the remainder term to depend on $\hat{K}^{(1)}$ so that, in general, the lowest order term in the effective Hamiltonian also appears to depend on $\hat{K}^{(1)}$. If we can find a $\hat{K}^{(1)}$ such that $\hat{R}=0$, $\hat{K}^{(1)}$ will not play a role in $\overline{\hat{H}^{(0)}_{\text{eff}}}$ but will instead appear as a gauge choice in Eq. (3). We will then recover the results presented in Sec. IV. A different condition, using the average of the remainder, could also be imposed: $\overline{\hat{R}}=0$. In the following, we choose the condition $\hat{R}=0$ instead of $\overline{\hat{R}}=0$ to avoid additional assumptions on pulse profiles.
We show two different solution methods for obtaining a vanishing remainder. The first is a trivial solution, $\hat{K}^{(1)}=0$ obtained by direct inspection of Eq. (52) where we see that all terms contain one $\hat{K}^{(1)}$ or its derivative. This proves that, for the simplest gauge choice, $\hat{K}^{(1)}=0$, Eq. (20) is correct.
We now establish a second, more general protocol to generate solutions to $\hat{R}=0$. We start by recalling that all observables are spanned by tensor products of Pauli matrices and the identity. $\hat{K}^{(1)}$ must then admit the form:
$\hat{K}^{(1)}=\sum_{\vec{\alpha}}c_{\vec{\alpha}}(t)\hat{\sigma}^{(\alpha_{1})}_{1}\cdots\hat{\sigma}^{(\alpha_{S})}_{S}=\sum_{\vec{\alpha}}c_{\vec{\alpha}}(t)\prod_{j}\hat{\sigma}^{(\alpha_{j})}_{j}$ (53)
where $\vec{\alpha}=\alpha_{1},...,\alpha_{S}$ and $S$ is the number of spins in the system. If the $c_{\vec{\alpha}}(t)$ are determined, then so is $\hat{K}^{(1)}$. We now substitute Eqs. (49) and  (53) into Eq. (52) and set $\hat{R}=0$. To do this, we will need to sum over a permutation of the locations of the product of Pauli operators appearing in (53). So let $\mathcal{P}$ denote a permutation of the sequence of S1S-1 zeros and $1$ one, then we rewrite terms in Eq. (52). For the first term in Eq. (52) we can use the following:
[…]
Whenever we consider a non-trivial permutation, the operator appearing just to the right of it is also permuted. With this substitution, Eq. (52) becomes:
[…]
where we have dropped the time dependence on $c_{\vec{\gamma}}$ for notational convenience. The expression contains long Pauli strings. To arrange them, we consider a permutation $\mathcal{P}_{0}$, so that the product $\prod_{l}\hat{\sigma}^{(\gamma_{l})}_{l}$ appears at the $s^{th}$ position in the nested commutator, and $1\leq s\leq p$. One can also show that:
[…]
$=\sum_{\lambda}\left((-1)^{s-1}(2i)^{p}\sum_{\delta_{1},...,\delta_{p-1}}\left(\prod_{r=1}^{p-1}\varepsilon^{\beta_{p+1-r}\delta_{r-1}\delta_{r}}\right)\varepsilon^{\beta_{1}\delta_{p-1}\lambda}\right)\left(\prod_{l\neq j}\hat{\sigma}^{(\gamma_{l})}_{l}\right)\hat{\sigma}^{(\lambda)}_{j}$, (56)
where we define $\delta_{0}=\alpha$ and $\beta_{s}=\gamma_{j}$ (which is valid since $1\leq s\leq p$ and the index on $\beta$ includes this range). Here $\varepsilon^{\alpha\beta\gamma}$ denotes the usual Levi-Civita symbol. We then use (56) to simplify (55). We exchange the sums over $\alpha,j,\vec{\gamma},\lambda$ with the sum over $p$ since they are independent, to obtain:
$\hat{R}=\sum_{\begin{subarray}{c}\alpha,j\\ \vec{\gamma},\lambda\end{subarray}}\left(A^{\alpha\lambda\vec{\gamma}}g^{\alpha}_{j}c_{\vec{\gamma}}-\frac{B^{\alpha\lambda\vec{\gamma}}}{\omega}\frac{dc_{\vec{\gamma}}}{dt}\right)\left(\prod_{l\neq j}\hat{\sigma}^{(\gamma_{l})}_{l}\right)\hat{\sigma}^{(\lambda)}_{j}$, (57)
where the coefficients are
$A^{\alpha\lambda\vec{\gamma}}$ $\equiv\sum_{p=1}^{\infty}\frac{i^{2p}}{p!}\frac{2^{p}p}{p+1}\sum_{\begin{subarray}{c}\beta_{1},...,\beta_{p-1}\\ k_{1},...,k_{p-1}\end{subarray}}G^{\beta_{1}}_{k_{1}}\cdots G^{\beta_{p-1}}_{k_{p-1}}\sum_{s=1}^{p}\left((-1)^{s-1}\sum_{\delta_{1},...,\delta_{p-1}}\left(\prod_{\begin{subarray}{c}r=2\\ r\neq p+1-s\end{subarray}}^{p-1}\varepsilon^{\beta_{p+1-r}\delta_{r-1}\delta_{r}}\right)\varepsilon^{\beta_{p}\alpha\delta_{1}}\varepsilon^{\gamma_{j}\delta_{p-s}\delta_{p+1-s}}\varepsilon^{\beta_{1}\delta_{p-1}\lambda}\right),$ (58)
$B^{\alpha\lambda\vec{\gamma}}$ $\equiv\sum_{p=1}^{\infty}\frac{i^{2p}(-2)^{p}}{(p+1)!}\sum_{\begin{subarray}{c}\beta_{1},...,\beta_{p}\\ k_{1},...,k_{p}\end{subarray}}G^{\beta_{1}}_{k_{1}}\cdots G^{\beta_{p}}_{k_{p}}\sum_{\delta_{1},...,\delta_{p-1}}\left(\prod_{r=2}^{p-1}\varepsilon^{\beta_{p+1-r}\delta_{r-1}\delta_{r}}\right)\varepsilon^{\gamma_{j}\alpha\delta_{1}}\varepsilon^{\beta_{1}\delta_{p-1}\lambda}.$ (59)
Since the generators are linearly independent, the condition $\hat{R}=0$ with the form in Eq. (57) gives as many equations as there are undetermined coefficients. These produce a set of coupled differential equations that can be solved generally to give a class of solutions for $\hat{K}^{(1)}$. That is, we must solve:
$A^{\alpha\lambda\vec{\gamma}}g^{\alpha}_{j}c_{\vec{\gamma}}-\frac{B^{\alpha\lambda\vec{\gamma}}}{\omega}\frac{dc_{\vec{\gamma}}}{dt}=0,\hskip 7.11317pt\text{all $\vec{\gamma}$}.$ (60)
There are solutions to Eqs. (60). The simplest solution is $c_{\vec{\gamma}}=0$, thus confirming that $\hat{K}^{(1)}=0$ is a valid gauge choice as discussed above. Different choices for $\hat{K}^{(1)}$ will, in general, require the construction of formal solutions to the set (60). In each case, the vanishing remainder term leads to:
$\overline{\hat{H}}_{\text{eff}}=\hat{H}_{0}+\sum_{p=1}^{\infty}\frac{i^{p}}{p!}\overline{[[\hat{K}^{(0)},\hat{H}_{0}]]_{p}}+\mathcal{O}(\omega^{-1}),$ (61)
which leads to Eq. (20). We note that $\overline{\hat{H}}_{\text{eff}}$ becomes time-independent only after time averaging. This is in contrast to methods using cancellation of individual terms without time averaging [Bukov2016].
@@ Appendix C Full Derivation of the Ising Interaction from the XXZ Hamiltonian
[…]
@@
[…]
@@
[…]
@@
[…]
@@
$\overline{\hat{H}_{\text{eff}}^{\text{S},(0)}}$ $=\hat{H}_{0}^{\text{S}}+\frac{1}{16}\left(\sum_{p=1}^{\infty}\frac{(-1)^{p}(4v)^{2p}}{(2p)!}\overline{G^{2p}_{1}}\right)D^{2}_{1}[\hat{H}_{0}^{\text{S}}]$
@@
$\overline{\hat{H}^{\text{S},(0)}_{\text{eff}}}=\sum_{j\neq j^{\prime}}\mathcal{V}_{j,j^{\prime}}\left[(J_{\perp}/2-4\delta JU^{(2)})\hat{\sigma}^{(1)}_{j}\hat{\sigma}^{(1)}_{j^{\prime}}+(J_{\perp}/2-4\delta JU^{(1)})\hat{\sigma}^{(2)}_{j}\hat{\sigma}^{(2)}_{j^{\prime}}\right]+\sum_{j\neq j^{\prime}}\mathcal{V}_{j,j^{\prime}}\left[\frac{J_{z}}{2}+4\delta J(U^{(1)}+U^{(2)})\right]\hat{\sigma}^{(3)}_{j}\hat{\sigma}^{(3)}_{j^{\prime}}$. (74)
$\overline{\hat{H}^{\text{S},(0)}_{\text{eff}}}$ allows us to use Floquet engineering to inter-relate a large class of starting Hamiltonians and effective Hamiltonians. For example, if we want to obtain $\overline{\hat{H}^{(0)}_{\text{eff}}}\propto\hat{H}_{ZZ}$, we need:
@@
$\overline{\hat{H}^{\text{S},(0)}_{\text{eff}}}=\frac{J_{z}+2J_{\perp}}{2}\sum_{j\neq j^{\prime}}\mathcal{V}_{j,j^{\prime}}\hat{\sigma}^{(3)}_{j}\hat{\sigma}^{(3)}_{j^{\prime}}$. (77)